OpenStudy (anonymous):

Write the quadratic equation in general quadratic form below. 3x^2 + 6x = 12

jimthompson5910 (jim_thompson5910):

This is similar to the one that I showed you before, so what would you do here to start?

OpenStudy (anonymous):

3x^2 + 6x - 12 = 0 \[x = (-b \pm \sqrt{b ^{2}-4ac})\div 2a\]

OpenStudy (anonymous):

the thing is I only have certain options so not all answers work even when I work the problem out..

jimthompson5910 (jim_thompson5910):

hmm not sure what to make of the pic, but all you're doing is subtracting 12 from both sides to go from 3x^2 + 6x = 12 to 3x^2 + 6x - 12 = 0 and that's it (because it's now in standard form) Unless you wanted to solve for x?

OpenStudy (anonymous):

yes, I have to solve for x

jimthompson5910 (jim_thompson5910):

oh ok, thanks for clarifying

OpenStudy (anonymous):

3x^2 + 6x = 12 3x^2 + 6x -12

jimthompson5910 (jim_thompson5910):

To solve for x, let's use the quadratic formula x = (-b+-sqrt(b^2-4ac))/(2a) x = (-(6)+-sqrt((6)^2-4(3)(-12)))/(2(3)) x = (-6+-sqrt(36-(-144)))/(6) x = (-6+-sqrt(180))/6 x = (-6+sqrt(180))/6 or x = (-6-sqrt(180))/6 x = (-6+6*sqrt(5))/6 or x = (-6-6*sqrt(5))/6 x = -1+sqrt(5) or x = -1-sqrt(5) So the solutions are x = -1+sqrt(5) or x = -1-sqrt(5)